OpenStudy (anonymous):

Few Core 2 questions

OpenStudy (anonymous):

Write \[\frac{ 1-x^2 }{ \sqrt[4]{x^3} }\] in the form \[x^p - x^q\]

OpenStudy (jdoe0001):

core2?

OpenStudy (anonymous):

Are you American? (No offense intended)

OpenStudy (jdoe0001):

guess so, as opposed to singaporean?

OpenStudy (anonymous):

Lol im canadian and never heard of that term tbh

OpenStudy (anonymous):

Either way it is part of the English AS maths course XD

OpenStudy (jdoe0001):

neither have I in math, that is I do know what core2 is in CSIT though

OpenStudy (anonymous):

\[(1-x^{2})(x^{-\frac{3}{4}})\]

OpenStudy (anonymous):

\[x^{-\frac{3}{4}}-(x^2*x^{-\frac{3}{4}})\]

OpenStudy (anonymous):

\[x^{-\frac{3}{4}}-(x^{2-\frac{3}{4}})\] \[x^{-\frac{3}{4}}-(x^{\frac {4}{4}-\frac{3}{4}})\] \[x^{-\frac{3}{4}}-x^{\frac {1}{4}}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

np